OpenStudy (anonymous):

write an expression for the situation the amount of money you spent if you started with $40 and now have d dollars

OpenStudy (anonymous):

okay so lets say that the variable S is the amount of money spent.

OpenStudy (anonymous):

okay so if you had 10 dollars and you bought a jacket and now your left with 5 . how much did the jacket cost

OpenStudy (anonymous):

5 dollars ?

OpenStudy (anonymous):

because 5 + 5 is 10

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

what about if you had 10 dollars and now your left with 2 dollars how much money did the jacket cost

OpenStudy (anonymous):

7 dollars

OpenStudy (anonymous):

hmm why do u say 7

OpenStudy (anonymous):

i meant 8

OpenStudy (anonymous):

becase 10 - 2 is 8

OpenStudy (anonymous):

because *

OpenStudy (anonymous):

good job! so you got that by doing 10 - 2 right? so if you had 40 dollars and now your left with d dollars how much did the thing you spent cost

OpenStudy (anonymous):

it would have to of cost 40 dollars 40 - 40 = 0

OpenStudy (anonymous):

hmm. well we dont know how much money we have left right nor do we have the cost of the thing that you bought

OpenStudy (anonymous):

all they tell us is that we are left with "d" dollars

OpenStudy (anonymous):

yeah so what would that be

OpenStudy (anonymous):

so lets say that S is the cost of the thing that we bought k? so S = 40 - d right?

OpenStudy (anonymous):

right so is that the answer?

OpenStudy (anonymous):

yup! do you understand why?

OpenStudy (anonymous):

s means spent right ?

OpenStudy (anonymous):

right it represents like the cost of the coat that we bought

OpenStudy (anonymous):

okay thank you ! can you help me with one more ?

OpenStudy (anonymous):

sure:)

OpenStudy (anonymous):

each persons share of baseball cards if 4 people share c cards equally

OpenStudy (anonymous):

very similar to the one we just did so EachPersonsShare = c/4